# Hyperparameter Optimization in Machine Learning

### v1

@@ 4.2 Acquisition Functions
A useful AF has to guide sampling so as to gain as much new information about high-performing hyperparameters as possible, given a limited sample $C$ obtained so far. In particular, it needs to balance the trade-off between exploration (i.e., sampling away from points in $C$ in order to increase coverage) and exploitation (i.e., refining the search by probing near current optima in $C$). We can gain intuition by considering two extremes. First, we could sample at our current best guess of where the minimum should be:
@@
A classic acquisition function, introduced in the seminal work by mockus1978application and frequently used in HPO libraries [snoek_practical_2012] and services [AMT2020], is the expected improvement (EI), defined as
@@
$\alpha_{\text{EI}}(\lambda|\beta,\hat{f}_{k})=\operatorname{\mathds{E}}\left[\max\left\{\beta-\hat{f}_{k}(\lambda),0\right\}\right].$
@@
Here, $\beta\in\mathbb{R}$ is a baseline, which is typically chosen as the minimum value of $f$ observed so far, $\beta=f(\hat{\lambda}_{k})$. EI has a simple closed-form expression in terms of $\mu_{k}(\lambda)$, $\sigma_{k}(\lambda)$, and in contrast to LCB does not come with a free parameter to be set.
A variant of this acquisition function is the probability of improvement [Kushner1964], given by
$\alpha_{\text{PI}}(\lambda|\beta,\hat{f}_{k})=\Phi(\nu(\lambda)),\text{ with }\nu(\lambda)=\frac{\beta-\mu_{k}(\lambda)}{\sigma_{k}(\lambda)}.$
@@
Another popular AF is known as Thompson sampling (TS) [Thompson1933, Hoffman2014]. The idea is to draw a sample path from our surrogate model and to search for its minimum point. Exact TS is intractable for a GP surrogate model, sample paths of which do not have a finite representation. In practice, we can draw a joint sample over a finite set of points333 Which costs $\mathcal{O}(b^{3})$ time and $\mathcal{O}(b^{2})$ memory, if $b$ is the size of the set. or approximate the GP posterior by a Bayesian linear model [Hernandez-Lobato2014].
@@
Other relevant families of AFs rely on information-theoretic criteria [Villemonteix2009, Hennig2012, Hernandez-Lobato2014, wang2017max]. While these have a sound theoretical foundation and have been shown to be more sample efficient, they typically come with higher computational cost and are more complex to implement than standard options such as EI or LCB. Given that the choice of the best AF can depend on the target function to optimize, one can also resort to a multi-armed bandit strategies to pick the most suitable AF from a portfolio of choices over the course of the optimization [hoffman2011portfolio, shahriari:2016:taking_human_out]. Techniques for maximizing acquisition functions include grid search, where the acquisition function is evaluated at a predefined set of points, and random search, which evaluates the acquisition function at randomly selected points. More sophisticated methods involve the use of gradient-based optimization, which leverages the gradient of the acquisition function to find local maxima more efficiently.
4.3 Tree Parzen Estimator
While Gaussian process based BO methods model the distribution $p(y|\lambda)$ of observations $y$ given configurations $\lambda$, the Tree Parzen Estimator (TPE) [bergstra2011algorithms, bergstra_making_2013] fits two conditional densities over $\Lambda$:
$l_{\beta}(\lambda)=p(\lambda|y<\beta),\quad g_{\beta}(\lambda)=p(\lambda|y\geq\beta).$
@@ 4.3 Tree Parzen Estimator
4.4 Decision Making Under Parallelism
@@
Training and validating a machine learning algorithm every time we evaluate the objective function might make HPO overly expensive. This is particularly true for modern machine learning algorithms applied on large datasets, such as deep neural networks, which often can take hours or days to train. In this chapter, we will consider how to obtain and exploit low-fidelity signals that serve as approximations of the objective function. These auxiliary metrics are much faster to evaluate than the full response function and often provide us with a sufficiently strong signal to guide the optimization process.
@@ 5.1 Multi-fidelity optimization
As noted in Section 2.5, an important aspect of manual search and hyperparameter tuning is fast turnaround time. To this end, practitioners will usually work on a subset of the training data in their first rough explorations. They will also not wait until training has converged, but rather check validation performance after a couple of iterations or epochs. Even though these early metrics are not entirely reliable, they are useful in order to home in on promising parts of the search space.
In multi-fidelity HPO [li2017hyperband, klein2017fast, swersky2014freeze], we assume that we have access to cheap-to-evaluate fidelities $f_{b}(\lambda)$ of the true objective function $f(\lambda)$, where the fidelity level is specified by the budget $b\in[b_{min},b_{max}]$. If we set the budget to the maximum budget $b=b_{max}$, we will recover the original objective function $f_{b_{max}}(\lambda)=f(\lambda)$, and we assume that the approximation of $f(\lambda)$ improves as $b$ increases. At the same time, we assume that the cost $c(\lambda,b)$ associated with evaluating $f_{b}(\lambda)$, for example the training time, increases with $b$ as well. While we can freely select $b$ during the optimization process, we are usually interested in the performance on the full budget $b_{max}$. More formally, we aim to find $\lambda^{*}\in\operatorname*{arg\,min}_{\lambda\in\Lambda}f_{b_{max}}(\lambda)$. In the next section, we will look at some commonly used fidelities that we can exploit to speed up the optimization process.
@@ 5.1.1 Iterations
For iterative machine learning algorithms, such as neural networks, we can interpret the number of iterations as budget and take the performance after each iteration as an approximation for the final performance [swersky2014freeze, li2017hyperband, jamieson2016non]. For neural networks we usually interpret epochs as iterations rather than single mini-batch updates, to reduce the overhead from validation error computations.
Figure 5.1 left shows the validation accuracy after each epoch of training for different hyperparameter configurations of a convolutional neural network. One can see that even after just a few epochs we are already able to visually distinguish between poor and well performing configurations. However, to identify the best configuration, we often have to evaluate the top performing configurations for the full number of epochs.
@@ 5.1.2 Training dataset subsets
Figure 5.1: Example learning curves of convolutional neural networks trained on CIFAR10.
The runtime of most machine learning algorithms scales linearly or even super-linearly with the number of training data points. Figure 5.1 right shows the validation error of a convolutional neural network trained on increasing subsets of the CIFAR10 dataset. Each color represents a different hyperparameter configuration. As one would expect, the validation error decays with the training dataset size. More interestingly, however, the ranking across hyperparameters remains relatively consistent across dataset subsets [bornschein-icml20]. This is related to recent work on scaling laws [kaplan-arxiv19, hoffmann-arxiv22], which show empirically that training loss of auto-regressive transformer [vaswani_attention_2017] models often follows a power law function with respect to the size of the training data.
In multi-fidelity optimization we can exploit this observation by interpreting the budget as the training dataset size [klein2017fast, swersky_multi-task_2013]. Thereby, to evaluate $f_{b}(\lambda)$, we only vary the training dataset size, but keep the validation set fixed, to avoid inducing additional noise. Most methods resample the training dataset uniformly at random for each evaluation of $f_{b}(\lambda)$, but also more sophisticated strategies, such as importance sampling, to select dataset subsets are possible [ariafar-aistats21]. Importantly, the class distribution needs to be preserved not to bias the performance estimation.
@@ 5.1.3 Cross-validation
Cross-validation (CV) is a popular technique to reduce the risk of overfitting on the validation dataset (see also Section 2.2.1). Rather than fixing a single training/validation fold, we use $K$ of them and average our target criterion over these. A main drawback of CV is its cost, which is $K$ times more expensive than to use a single fold. We can approximate it via multi-fidelity, by using the number of folds $b=1,\dots,K$ as budget.
Compared to the previous two fidelites, here we do not assume that the performance increases with the budget but rather that the variance estimate of the generalization performance decreases with the number of folds.
@@ 5.2 Methods based on random search
In this section, we will review algorithms which exploit multiple fidelities and suggest new configurations by random sampling. We will begin by clarifying the difference between synchronous and asynchronous decision making. Random-search based synchronous multi-fidelity methods have first been proposed in the multi-armed bandits (MAB) literature. In the following Section 5.3 we will concentrate on model-based multi-fidelity methods, where decision making is also driven by a surrogate model.
The methods discussed here assume that the evaluation costs $c(\lambda,b)$ for every fixed $\lambda$ are linear (or affine linear) in $b$.
5.2.1 Synchronous and asynchronous decision making
Multi-fidelity methods can be classified into synchronous and asynchronous. In the former, decisions about stopping or promoting trials are delayed until a certain batch of evaluations has finished. In successive halving or Hyperband, to be discussed shortly, all evaluations in a rung have to finish before a fraction of these trials are allowed to continue. In contrast, a decision whether to stop or continue a trial must be done instantaneously in asynchronous methods, based on the data available at that time. ASHA is an example for this type of scheduling.
Synchronous multi-fidelity methods are simpler to understand and analyze. On the other hand, they are harder to implement efficiently in an asynchronously parallel computation environment, and even sophisticated implementations tend to clock more idle time.
@@ 5.2.2 Successive halving
Successive Halving (SH) [karnin-icml13, jamieson2016non] is a MAB strategy that aims to identify the best hyperparameter configuration out of a finite set of configurations. Given the maximum and minimum budget $b_{max}$ and $b_{min}$ for evaluating a single hyperparameter configuration, we first define a set of rung levels
=$\mathcal{R}=\{b_{min}\eta^{k}\;|\;k=0,\dots,K\}.$
We assume, for simplicity, that $b_{max}/b_{min}=\eta^{K}$, where $K\in\mathbb{N}$. Here $\eta\geq 2$ defines the halving constant and determines how aggressively successive halving will prune configurations. We start by drawing $n$ hyperparameter configurations uniformly at random and evaluate them on the first rung level, i.e each configuration is evaluated with a budget of $b_{min}$. For convenience the number of initial configurations is often set to $n=\eta^{K}$ but this can be also arbitrary. Afterwards, successive halving promotes $\frac{1}{\eta}$ of the configurations, sorted by their validation performance to the next higher rung level $b_{min}\eta$. These steps are iterated until we reach the highest rung with level $b_{max}$. See Figure 5.2 for an visualization.
While conceptually simple, successive halving showed strong performance in the past and even out-performed more sophisticated Bayesian optimization techniques that only operate on the full budget [jamieson2016non, li2017hyperband]. However, its performance relies on a high correlation between the ranking of hyperparameter configurations across rung levels. In other words, the performance of a configuration on the first rung level should be representative of its performance on the highest rung level. Otherwise configurations that perform best on $b_{max}$ will be discarded too quickly.
Figure 5.2: Successive halving splits the maximum budget $b_{max}$ into a set of exponentially increasing rungs. Then, in each iteration, successive halving evaluates all configurations on the current rung and promotes the top $\frac{1}{eta}$ of the configurations to the next rung.
How is parallel computation exploited by SH? Recall the notation of synchronous versus asynchronous parallelization from Section 2.4, and assume that $W$ workers are available. For synchronous scheduling, we can start $\min(W,k)$ jobs in parallel, where $k$ is the number of free slots remaining in the currently processed rung. Here, $W-k$ workers remain idle, which is particularly wasteful for the higher rungs. For example, if $n=\eta^{K}$, the highest rung has a single slot only, so $W-1$ workers will be idle for the maximum budget $b_{max}$. The same issue arises for asynchronous scheduling. A remedy will be discussed in Section 5.2.4.

### v2

@@ 4.2.3 Acquisition functions
A useful AF has to guide the choice of the next candidate so as to gain as much new information about high-performing hyperparameters as possible, given a limited sample $\mathcal{D}$ obtained so far. In particular, it needs to balance the trade-off between exploration (i.e., looking away from points in $\mathcal{D}$ in order to increase coverage) and exploitation (i.e., refining the search by probing near current optima in $\mathcal{D}$). We can gain intuition by considering two extremes. First, we could sample at our current best guess of where the minimum should be:
@@
A classic AF, introduced in the seminal work by mockus1978application and frequently used in HPO libraries [snoek_practical_2012] and services [AMT2020], is the expected improvement (EI), defined as
@@
$\alpha_{\text{EI}}(\lambda|\beta,\hat{f}_{k})\doteq\operatorname{\mathds{E}}\left[\max\left\{\beta-\hat{f}_{k}(\lambda),0\right\}\right].$
@@
Here, $\beta\in\mathbb{R}$ is a baseline, which is typically chosen as the minimum value of $f$ observed so far, $\beta=f(\hat{\lambda}_{k})$. In contrast to LCB, EI does not come with a free parameter to be set. In addition, it can be shown to have a simple closed-form expression in terms of $\mu_{k}(\lambda)$ and $\sigma_{k}(\lambda)$:
$\alpha_{\text{EI}}(\lambda|\beta,\hat{f}_{k})=\sigma_{k}(\lambda)\nu(\lambda)\Phi\left(\nu(\lambda)\right)+\sigma_{k}(\lambda)\phi\left(\nu(\lambda)\right)$ (4.14)
where
$\nu(\lambda)\doteq\frac{\beta-\mu_{k}(\lambda)}{\sigma_{k}(\lambda)}$
and where $\Phi$ and $\phi$ denote the cumulative distribution function and the probability density function of the standard Gaussian distribution, respectively.
Equations (4.13) where $\alpha$ is the expected improvement and (4.12) specify a particular simple instantiation of a GP-based BO algorithm that may be used for HPO. As an illustrative example, two steps of the Bayesian optimization algorithm using EI as an AF are shown in Figure 4.1.
Figure 4.1: Left: Suppose we have measured the response function (plotted in red) on a small set of points. Fitting a Gaussian process on these points yields a mean function (plotted in black) with and a variance function ($\pm$ one standard deviation is filled in gray). If $\lambda$ falls between 7 and 11, the uncertainty is very high, which yields the right bump in the expected improvement (plotted in green), allowing us to explore a region where potentially a smallest response might be found. The left bump of the expected improvement falls in an exploitation region, which in this example looses against exploration. Hence, the next candidate is $\lambda=8.5$. Right: once the response is measured at the new candidate, the Gaussian process is trained again. Now uncertainty shrinks significantly and the new expected improvement has just one bump around 2.7, which is already very close to the minimum of the true response. This example takes inspiration from Figure 11 in [Jones1998:Efficient].
Other acquisition functions.
An earlier and simpler variant of EI is the probability of improvement [Kushner1964], given by
$\alpha_{\text{PI}}(\lambda|\beta,\hat{f}_{k})\doteq P\left[\hat{f}_{k}(\lambda)<\beta\right]=\Phi(\nu(\lambda)).$ (4.15)
@@
Another popular AF is known as Thompson sampling (TS) [Thompson1933, Hoffman2014]. The idea is to draw a sample path from our surrogate model and to search for its minimum point. Exact TS is intractable for a GP surrogate model, sample paths of which do not have a finite representation. In practice, we can draw a joint sample over a finite set of points333 Which costs $\mathcal{O}(b^{3})$ time and $\mathcal{O}(b^{2})$ memory, if $b$ is the size of the set. [unify notation please] or approximate the GP posterior by a Bayesian linear model [Hernandez-Lobato2014].
@@
Other relevant families of AFs rely on information-theoretic criteria [Villemonteix2009, Hennig2012, Hernandez-Lobato2014, wang2017max]. While these have a sound theoretical foundation and have been shown to be more sample efficient, they typically come with higher computational cost and are more complex to implement than standard options such as EI or LCB. Given that the choice of the best AF can depend on the target function to optimize, one can also resort to a multi-armed bandit strategies to pick the most suitable AF from a portfolio of choices over the course of the optimization [hoffman2011portfolio, shahriari:2016:taking_human_out].
Optimizing acquisition function in practice
Techniques for maximizing AFs include grid search, where the AF is evaluated at a predefined set of points, and random search, which evaluates the AF at randomly selected points. More sophisticated methods involve the use of gradient-based optimization, which leverages the gradient of the AF to find local maxima more efficiently.
4.2.4 Aleatoric and epistemic uncertainty
In (4.4) we have introduced a noise term $\epsilon$ in order to account for the stochastic nature of the learning algorithm itself. This derives from the internal randomness of the training algorithm, (for example due to initial weights when training a deep network). In general, the amount of this aleatoric uncertainty may depend on the current hyperparameters and should not be confused with the epistemic uncertainty which arises because we may have not (yet) evaluated a large enough set of hyperparameter vectors. The latter uncertainty is crucial to allow explorations in regions not yet assessed, while the former should not mislead us to acquire points where uncertainty cannot be reduced.
A heteroscedastic approach can help disentangle these types of uncertainties, as proposed by [griffiths:2021]. In their approach, distinct Gaussian processes are trained aiming to predict the unknown response and the noise separately.
4.2.5 Tree Parzen estimator
The Tree Parzen Estimator (TPE) is a probabilistic model to efficiently explore complex search spaces. It models the objective function as a pair of conditional probability distributions: one for configurations that yield better results than a threshold and another for those that perform worse. By iteratively selecting hyperparameter values that maximize the expected improvement, TPE focuses the search on promising regions. This approach is particularly well-suited for high-dimensional, non-convex optimization problems and balances exploration and exploitation to accelerate convergence toward optimal configurations. While Gaussian process based BO methods model the distribution $p(y|\lambda)$ of observations $y$ given configurations $\lambda$, the Tree Parzen Estimator (TPE) [bergstra2011algorithms, bergstra_making_2013] fits two conditional densities over $\Lambda$:
$l_{\beta}(\lambda)=p(\lambda|y<\beta),\quad g_{\beta}(\lambda)=p(\lambda|y\geq\beta).$ (4.16)
@@ 4.2.5 Tree Parzen estimator
4.3 Other approaches to model-based HPO
Not all model-based methods follow a Bayesian optimization framework. ilievski2017efficient use a deterministic radial basis function model for $\hat{f}$ which they interpolate on observations of $f$, while hazan2017hyperparameter focus on Boolean hyperparameters and use a sparse Fourier basis model for $\hat{f}$, in conjunction with heuristics to estimate the most sensitive hyperparameter entries. They then use a base global optimizer to perform HPO on a restriction of the original hyperparameter space. The works by lorraine2018stochastic and mackay2019self, linked to gradient-based HPO may also be interpreted as an attempt to learn a surrogate model of the entire learning algorithm (applied to a specific dataset), whereby the output space of $\hat{f}$ coincides with the hypothesis space of $A$ itself, that is $\hat{\mathcal{Y}}=\mathcal{H}$ in (4.1). We cover gradient-based HPO in Chapter 7.
Furthermore, not all methods follow the iterative schema of Eq. (4.3). Another possibility is to first learn $\hat{f}$ and then use it to suggest one or several hyperparameter configurations in a second stage, following a parallel scheme. Examples of this latter approach include [feurer2015initializing, fusi2018probabilistic]. These methods, however, assume the presence of multiple tasks and are closer in nature to meta or transfer learning settings.
4.4 Example: model-based vs random search
We illustrate the difference between model-based and random search in a concrete example of tuning regularization parameters of an ML model to prevent overfitting. We tuned the L1 and L2 regularization terms, respectively alpha and lambda, of the XGBoost algorithm on the UCI direct marketing binary classification dataset from Kaggle. The goal was to minimize the AUC. Figure 4.2 presents a comparison between random search and BO (model-based), which models the objective function through a GP and uses the expected improvement as the acquisition function.
Each experiment was conducted using 50 different random seeds, and the results were reported as the average and standard deviation. The left and middle plots reveal that BO consistently suggests better-performing hyperparameters than random search. The right plot further demonstrates that BO consistently outperforms random search across all evaluated hyperparameter configurations.
While BO is particularly effective in a sequential setting, random search remains a valuable approach, especially in distributed environments where many configurations can be evaluated in parallel, significantly reducing wall-clock time.
Figure 4.2: Top: Hyperparameters suggested by random search (left) and BO (right). The $x$-axis and $y$-axis are the hyperparameter values for alpha and lambda and the color of the points represent the quality of the hyperparameters (yellow: better AUC; violet: worse AUC). Bottom: Best model score obtained so far ($y$-axis, where lower is better) as more hyperparameter evaluations are performed ($x$-axis).
4.5 Discussion
Model-based methods tend to be somewhat more complex than other HPO techniques, both conceptually and from an implementation standpoint. They require fitting a (probabilistic) surrogate model with relatively few function evaluations, and to update this fit online whenever new observations become available. Moreover, the optimal decision on where to evaluate $f$ next needs to be approximated via an AF depending on the surrogate model, as we saw in Section 4.2.3. Thus, while there is a thriving ecosystem of free academic software, it is not rare that model-based HPO techniques are implemented and proposed as commercial products [sigopt-web-page, golovin2017google, AMT2020]. The development of MB techniques requires several design choices to be made, such as the concrete representation of $\hat{f}$. Some configurations may be left to be set by the final user. While authors often provide “default values”, the presence of many such parameters may make some model-based techniques less accessible. See, e.g., the list of required inputs for the algorithms proposed by hazan2017hyperparameter or falkner2018bohb.
While the foremost representatives of these type of methods – tied to BO – are classically black-box procedures, model-based techniques may greatly benefit from direct access to the underlying learning algorithm. For instance, several works [swersky2014freeze, klein2017fast, falkner2018bohb] use information gathered from partial execution of $A$ to stop unpromising evaluations, which may be additionally resumed at a later stage [swersky2014freeze] as we will discuss in more details in the next chapter. Before doing so, we spend some words on what is possibly the biggest practical limitation of (naive) model-base methods: parallelizability.
4.5.1 Decision making under parallelism
@@
One of the main challenges of HPO is that training and validating a machine learning algorithm each time the objective function is evaluated can be expensive. This is particularly true for modern machine learning algorithms applied to large datasets, such as deep neural networks, which can often take hours or days to train. In this chapter, we consider how to obtain and exploit low-fidelity signals that serve as approximations of the objective function. These auxiliary metrics are much faster to evaluate than computing the full objective function and often provide a strong enough signal to guide the optimization process.
We begin with an overview of different types of fidelities in Section 5.1. Section 5.2 explores methods for multi-fidelity optimization based on randomly sampling configurations. Finally, in Section 5.3, we examine multi-fidelity approaches that leverage probabilistic models to guide the search.
@@ 5.1 Multi-fidelity optimization
As noted in Section 2.5, an important aspect of manual search and hyperparameter tuning is fast turnaround time. To achieve this, practitioners typically work with a subset of the training data during their initial explorations. They also do not wait for training to fully converge but instead check validation performance after a few iterations or epochs. While these early metrics are not entirely reliable, they are useful for identifying promising regions of the search space.
In this section, we discuss how these manual practices can be transformed into algorithmic tools that enable automated early stopping of underperforming runs.
Multi-fidelity HPO [li2017hyperband, klein2017fast, swersky2014freeze] extends the HPO formulation from Section 2.2 by taking the budget that we spend to train a hyperparameter configuration into explicit account. We assume access to inexpensive approximations of the true objective function $f(\lambda)$, denoted as $f_{b}(\lambda)$, where the fidelity level is determined by the budget $b\in[b_{min},b_{max}]$.
If we set the budget to its maximum value, $b=b_{max}$, we recover the original objective function, $f_{b_{max}}(\lambda)=f(\lambda)$. From this perspective, standard HPO can be seen as a special case of multi-fidelity HPO. Additionally, we assume that the approximation of $f(\lambda)$ improves as $b$ increases. However, the cost $c(\lambda,b)$ associated with evaluating $f_{b}(\lambda)$—such as training time—also increases with $b$.
While we can freely choose $b$ during the optimization process, our primary interest lies in the performance at the full budget $b_{max}$. More formally, our goal is (still) to find $\lambda^{*}\in\operatorname*{arg\,min}_{\lambda\in\Lambda}f_{b_{max}}(\lambda)$, which coincides with the formalization of Chapter 2.
In the next section, we explore commonly used fidelity measures that can be leveraged to accelerate the optimization process.
@@ 5.1.1 Iterations
For iterative machine learning algorithms, such as neural networks, we can interpret the number of iterations as the budget and use the performance at each iteration as an approximation of the final performance [swersky2014freeze, li2017hyperband, jamieson2016non]. In the case of neural networks, we typically consider epochs as iterations rather than individual mini-batch updates to reduce the overhead of validation error computations.
Figure 5.1 (left) illustrates the validation loss after each epoch of training for different hyperparameter configurations of a multi-layer perceptron. Even after just a few epochs, we can visually distinguish between poorly performing and well-performing configurations. However, to reliably identify the best configuration, it is often necessary to evaluate the top-performing candidates for the full number of epochs.
@@ 5.1.2 Training dataset subsets
Figure 5.1: Left: Validation loss of different neural networks evaluated after each epoch of training. Right: Validation error of a support vector machine trained on different subsets of the training data. Each color represents a different hyperparameter configuration.
The runtime of most machine learning algorithms scales linearly or even super-linearly with the number of training data points. Figure 5.1 (right) shows the validation error of a support vector machine trained on increasing subsets of the MNIST dataset, with each color representing a different hyperparameter configuration. As expected, the validation error decreases as the training dataset size increases. More interestingly, however, the ranking of hyperparameter configurations remains relatively consistent across dataset subsets [bornschein-icml20].
In the context of deep neural networks, this observation is related to recent research on scaling laws [kaplan-arxiv19, hoffmann-arxiv22], which empirically demonstrate that the training loss of autoregressive transformer models [vaswani_attention_2017] often follows a power-law function with respect to the size of the training dataset.
In multi-fidelity optimization, we can exploit this phenomenon by interpreting the budget as the training dataset size [klein2017fast, swersky_multi-task_2013]. To evaluate $f_{b}(\lambda)$, we vary only the training dataset size while keeping the validation set fixed to avoid introducing additional noise. Most methods resample the training dataset uniformly at random for each evaluation of $f_{b}(\lambda)$, though more sophisticated strategies, such as importance sampling, can also be used to select dataset subsets [ariafar-aistats21]. Importantly, global statistics such as the class distribution must be preserved to avoid biasing performance estimation.
@@ 5.1.3 Cross-validation
As discussed in Section 2.2.1, cross-validation (CV) is a widely used technique for estimating the generalization error of a machine learning method and reducing the risk of overfitting. Rather than relying on a single training/validation split, we divide the dataset into $K$ subsets or folds, using each fold for validation in turn and averaging our target criterion across them. The main drawback of CV is its computational cost, which is $K$ times higher than using a single fold. However, we can approximate it using a multi-fidelity approach by treating the number of folds, $b=1,\dots,K$, as the budget.
Unlike the previous two fidelities, where performance typically improves with increasing budget, CV does not necessarily exhibit this behavior. Instead, the key assumption is that the variance of the generalization performance estimate decreases as the number of folds increases.
@@ 5.2 Methods based on random search
In this section, we review algorithms that leverage multiple fidelities and propose new configurations through random sampling. Random search based multi-fidelity methods were first introduced in the multi-armed bandit (MAB) literature. We discuss how a simple MAB approach can dynamically halt the evaluation of poorly performing configurations. Additionally, we discuss how these methods can be adapted for efficient execution in large-scale distributed settings. The methods covered here assume that the evaluation costs, $c(\lambda,b)$, for any fixed $\lambda$ are linear (or affine linear) in $b$.
@@ 5.2.1 Successive halving
Successive halving (SH) [karnin-icml13, jamieson2016non] is an MAB strategy that aims to identify the best hyperparameter configuration from a finite set of configurations. Given the maximum and minimum budgets, $b_{max}$ and $b_{min}$, for evaluating a single hyperparameter configuration, we first define a set of rung levels:
=$\mathcal{R}=\{b_{min}\eta^{k}\mid k\in[K]\}$
We assume, for simplicity, that $b_{max}/b_{min}=\eta^{K}$, where $K\in\mathbb{N}$. Here, $\eta\geq 2$ defines the halving constant and determines how aggressively SH prunes configurations.
Algorithm 2 Successive-Halving-HPO
1:Initial budget $b_{min}$, maximum budget $b_{max}$
2:Set of $n$ candidate configurations $\mathcal{C}_{0}=\{\lambda_{1},\lambda_{2},\dots\lambda_{n}\}$
3:Halving constant $\eta\geq 2$
4:for $i\in[|\mathcal{R}|]$ do
5:     $\{f_{r_{i}}(\lambda):\lambda\in\mathcal{C}_{i-1}\}$triangleright Evaluate configurations on the rung
6:     $k\leftarrow\lfloor|\mathcal{C}_{i-1}|/\eta)\rfloor$triangleright Number of configurations to keep
7:     $\mathcal{C}_{i}\leftarrow\underset{\lambda\in\mathcal{C}_{i-1}}{\arg\mathrm{top}\text{-}k}f_{r_{i}}(\lambda)$triangleright Keep $k$ best performing configurations
8:end for
Algorithm 2 shows pseudo code for SH. Given a set $\mathcal{C}_{0}$ of $n$ hyperparameter configurations, usually drawn uniformly at random, we evaluate them at the first rung level, i.e., each configuration is evaluated with a budget of $b_{min}$ (Line 2). For convenience, the number of initial configurations is often set to $n=\eta^{K}$, but this can also be arbitrary. Afterwards, SH promotes $\frac{1}{\eta}$ of the configurations, sorted by their validation performance, to the next higher rung level $b_{min}\eta$ (Line 4). These steps are iterated until we reach the highest rung level, $b_{max}$. See Figure 5.2 for a visualization.
While conceptually simple, SH has shown strong performance in the literature and even outperformed more sophisticated Bayesian optimization techniques that only operate on the full budget [jamieson2016non, li2017hyperband]. However, its performance relies on a high correlation between the ranking of hyperparameter configurations across rung levels. In other words, the performance of a configuration at the first rung level should be representative of its performance at the highest rung level. Otherwise, configurations that perform best at $b_{max}$ may be discarded too quickly.
Figure 5.2: Successive halving splits the maximum budget $b_{max}$ into a set of exponentially increasing rungs. Then, in each iteration, successive halving evaluates all configurations on the current rung and promotes the top $\frac{1}{\eta}$ of the configurations to the next rung.